OpenStudy (anonymous):

SOMEONE HELP IM DYING MY SCHOOL WILL NOT LET ME BECOME A STUDENT AT THIER SCHOOL UNLESS I FINISH THIS CLASS!! ALGEBRA II 2. Adam and Beth are visiting an amusement park and have decided to ride the carousel. Adam picked a horse on the outside edge and Beth chose a dragon on the inside, closer to the center. Part 1: Do Adam and Beth travel the same distance during the ride? Choose a distance that each seat (horse and dragon) sits from the center and use the radius to determine how far each would travel during one rotation.

OpenStudy (anonymous):

So far I have Adam r = 8 Beth r = 4 Then I used the formula C = 2pir 2*3.14*8 = 50.24 2*3.14*4 = 25.12

OpenStudy (anonymous):

@mathmale @ganeshie8 !!!!! I love you for even trying!! PLease !!

OpenStudy (mathmale):

Dear nsmohammad, I love you for trying too. Your formula C=2pi*r is for finding the circumference of an entire circle. This circumference depends directly upon r. If r increases, so does the circumference. So if the 2 kids are different distances from the center of the merry-go-round, will they travel the same distance or different distances as the wheel turns?

OpenStudy (mathmale):

You've already done the work. Just go back to the two equations you typed out and compare the results. Same or different? How would you answer this question in words?

OpenStudy (anonymous):

The results are different. Adam's position has a longer radius, so he gets more out of the ride. Beth's position has a shorter radius so she will not have as much fun. Im sorry I am not very good at this

OpenStudy (mathmale):

Wonderful choice of words! "Gets more out of the ride" / travels further because he is further away from the center (his radius is greater. The other person, Beth, is closer to the center, so her radius is smaller, and "she will not have as much fun" / she will travel a shorter distance than will Adam.

OpenStudy (mathmale):

Your language is poetic and I like it! But you have to mention DISTANCE in each case. Adam travels ... can you finish this sentence, using that word?

OpenStudy (anonymous):

Adam travels a longer distance than Beth. I am honestly not good at math where you have to write only. I am better in algebra 2 in things like solving equations, logarithms, etc. I did not expect to do statistics You have to help me out here

OpenStudy (mathmale):

Although it's great to be good at algebra, at solving equations, logs, etc., it's quite important to know how to explain your results in words. come on, nsm, stop saying "i'm not good at math." What you don't know, you can learn, and you're already demonstrating this.

OpenStudy (mathmale):

Quick: Let's move on to the next problem you want to discuss.

OpenStudy (anonymous):

Ok there is a part 2

OpenStudy (anonymous):

Part 2: Choose an angle of rotation. Using complete sentences, compare the distance Adam and Beth will travel during this angle measurement.

OpenStudy (mathmale):

Type that out for me, please.

OpenStudy (anonymous):

I have to choose an angle then convert it to radians and use the arc length formula to find the distance they traveled during the angle measurement

OpenStudy (mathmale):

The "arc length formula" reads like this:\[s=r* \theta\]

OpenStudy (mathmale):

Note that that angle "theta" MUST be expressed in radians. What kind of help do you need to be able to solve this problem?

OpenStudy (anonymous):

Does it matter which angle I choose? I know how to convert degrees to radians and radians to degrees. I just don't know if it matters what angle I choose.

OpenStudy (mathmale):

Just supposing that the central angle were pi/2 radians (or 90 degrees), and that the radius is 3 meters. The distance traveled is the arc length, which, in turn, is (3 meters)(pi/2 radians).

OpenStudy (mathmale):

I think you just choose whatever angle you want. for simplicity, choose pi/3 rad (60 deg). How far will the person on the outside travel if that p erson is 3 m. from the center?

OpenStudy (mathmale):

If r = radius = 3 meters, and if theta = the central angle, in radians = pi/3, Then the arc length (which represents the distance that person will travel in the arc) is \[s=r*\theta \]

OpenStudy (mathmale):

Go ahead and multiply to find the arc length / distance traveled.

OpenStudy (anonymous):

3.14?

OpenStudy (mathmale):

Let's review this: the radius, r, is 3 meters; the central angle, theta, is pi/3 radians. You've multiplied r by the central angle in radians and have obtained pi, or approx. 3.14 meters. Right. That's how far this person 3 meters from the center has traveled in an arc when the central angle is pi/3 radians.

OpenStudy (anonymous):

Would the other person have a longer radius?

OpenStudy (mathmale):

You have done that problem correctly. Just be sure to write 3.14 meters, not just 3.14. I invented that 3 meter radius for our use as an example. If we decide that the other person is 4 meters from the center, then the distance traveled by that person (in an arc) is s = (4 m)(pi/3 rad).

OpenStudy (anonymous):

So for my radiuses 4 and 8 the answer if I was using 90 degrees it would be S = 4* 3.14/2 = 6.28 S = 8*3.14/2 = 12.56

OpenStudy (mathmale):

yes, except you must type in the units of measurement (inches, feet, meters, or whatever).

OpenStudy (anonymous):

I am sorry I am keeping you up so late I have to finish 13 assignments in 3 days or else my school willn not accept me

OpenStudy (mathmale):

Although I'd like to stop now, I'd be delighted to continue working with you later. I'd suggest you post y our questions and see if any others are able to help; if you are at all dissatisfied, just send me a message and I'll help you with those problems as soon as possible.

OpenStudy (anonymous):

So was my work right?

OpenStudy (anonymous):

One more question please I have my answer but I know it's wrong

OpenStudy (mathmale):

Have you posted it? Please do that separately from the problem we've just finished.